How do you define Chemical potential of a species i in U, H, G and A. and how they are
related to each other?
Answer #1

Chemical potential is energy required to add or release a particle from a given species.

So In tyermodynamic relations change in internal energy of a recation we write as

dU=TdS-PdV+\Sigma_i^N\mu_i dN_i

So chemical potential gor i th particle change on constant Volume ans entroy we write as

\mu_i =\frac{\partial U}{\partial N_i}_{S,V,N_{i\neq j}}

Now gibbs free energy

G=U + PV - TS

so,

dG = dU + PDV + V DP - TIS - SDT

now substituting for dU

we get

dG = -SDT +VdP + MidN,

so,we define for constant T and P

\mu_i =\frac{\partial G}{\partial N_i}_{T,P,N_{i\neq j}}

similarly

H = U + PV

we write

dH=dU+PdV+VdP

Substitute for dU

dH=TdS+VdP+\Sigma _i^N\mu_i dN_i

so,

\mu_i =\frac{\partial H}{\partial N_i}_{S,P,N_{i\neq j}}

simlarly

A= U-TS

So,

dA= dU-TdS-SdT

so,

dA=-PdV-SdT+\Sigma_i^N\mu_idN_i

so,

\mu_i =\frac{\partial A}{\partial N_i}_{V,T,N_{i\neq j}}